40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 26 22:24:28 2014, Jenne, Update, LSC, DARM loop improved, other work 8x
    Reply  Mon Dec 1 17:43:22 2014, Jenne, Update, LSC, Reset Yarm trans normalization 
Message ID: 10737     Entry time: Wed Nov 26 22:24:28 2014     Reply to this: 10743
Author: Jenne 
Type: Update 
Category: LSC 
Subject: DARM loop improved, other work 

[Jenne, Koji]

We have done several things this evening, which have incrementally helped the lock stability.  We are still locking CARM and DARM on ALS, and PRMI on REFL165.

  • Saw peaks in CARM error signal at 24Hz and 29 Hz, so put in moderate-Q resonant gains. 
  • DARM at low frequency was much noisier than CARM.  We discovered that we had put in a nice boost at some point for CARM in FM1, but hadn't transferred that over to DARM.  Copying FM1 from CARM to DARM (so replacing an integrator with a boost below ~10Hz) dropped the DARM noise down to match the CARM noise at low frequencies.
  • Koji noticed that we were really only illuminating one quadrant of the Xend QPD, so we aligned both trans QPDs.  Also, I reset the transmission normalization so that all 4 diodes (Thorlabs and QPDs at each end) all read 1 with good alignment.
  • We've got some concerns about the ASS.  It needs some attention and tuning.
    • The X ASS needs an overall gain of about 0.3.  This may be because I forgot to put the new lower gains into the burt restore after Rana's oplev work, or this may be something new.
    • When Koji did a very careful arm alignment, we turned on the Y ASS and saw it methodically reduce the transmitted power.  Mostly it was moving ETMY in yaw.  Why is the DC response of the ASS not good?  The oplev work shouldn't have affected DC.
    • We don't like the way the ASS offloads the alignment.  Maybe there's a better way to do it overall, but one thought is to have an option to offload (for long-term alignment fixing, so maybe once a day) and another option to just freeze the current output (for the continual tweak-ups that we do throughout the evening).  We'd want the ASS to start up again with these frozen values, and not clear them.
  • ETMY was being fussy, in the same way that ETMX had been for the last few months.  I went down to squish the cables, and found that it was totally not strain-relieved, and that the cable was pulling on the connector.  I have zip tied the cable to the rack so that it's not pulling anymore.
  • At high arm powers, it is hard to see what is going on at the AS port because there is so much light.  Koji has added an ND filter to the AS camera so that we can more easily tweak alignment to improve the contrast.

Something that has been bothering me the last few days is that early in the evening, I would be able to get to very high arm powers, but later on I couldn't.  I think this has to do with setting the contrast at the AS port separately for the sideband versus the carrier.  I had been minimizing the AS port power with the arms held off resonance, PRMI locked.  But, this is mostly sideband.  If instead I optimize the Michelson fringes when the arms are held with ALS at arm powers of 1, and PRM is still misaligned, I end up with much higher arm powers later.  Some notes about this though:  most of this alignment was done with the arm cavity mirrors, specifically the ETMs, to get the nice Michelson fringes.  When the PRM is restored and the PRMI locked, the AS port contrast doesn't look very good.  However, when I leave the alignment alone at this point, I get up to arm powers above 100, whereas if I touch the BS, I have trouble getting above 50.


Around GPS time 1101094920, I moved the DARM offset after optimizing the CARM offset.  We were able to see a pretty nice zero crossing in AS55, although that wasn't at the same place as the ALS diff zero offset (close though).  At this time, the arm powers got above 250, and TRY claimed almost 200.  These are the plots below, first as a wide-view, then zoomed in.  During this time, PRCL still has a broadband increase in noise when the arm powers are high, and CARM is seeing a resonance at a few tens of Hz.  But, we can nicely see the zerocrossing in AS55, so I think there's hope of being able to transition DARM over. 

DARMcrossing_Power.png

DARMcrossing_ErrCtrl.png

DARMcrossing_AuxErr.png

DARMcrossing_Angles.png

Now, the same data, but zoomed in more.

DARMcrossing_Power_Zoom.png

DARMcrossing_ErrCtrl_Zoom.png

DARMcrossing_AuxErr_Zoom.png

DARMcrossing_Angles_Zoom.png


During the 40m meeting, we had a few ideas of directions to pursue for locking:

  • Look into using POX or POY as a proxy for POP and instead of REFL, for CARM control.  Maybe we have some nice juicy SNR.
  • Check the linearity of our REFL signals by holding the arms on (or close to) resonance, then do a swept sine exciting CARM ctrl and taking a transfer function to the RF signals.
  • Q is going to look into the TRX QPD, since he thought it looked funny last week, although this may no longer be necessary after we put the beam at the center of the QPD.
  • Koji had a thought for making it easier to blend the CARM error signals.  What if we put a pole into the ALS CARM signals at the place where the final coupled cavity pole will be, and then compensate for this in the CARM loop.  Since any IR signals will naturally have this pole, we want the CARM loop to be stable when it's present.
  • Diego tells us that the Xarm IR beatnote is basically ready to go.  We need to see how big the peak is so we can put it into the frequency counter and read it out via EPICS.  The freq counter wants at least -15dBm, so we may need an amplifier.
ELOG V3.1.3-